17
Sep
2013
08:13

Update: Normal operation has now been restored. Should customers still have any problems then please contact us immediately. A technical note update will be posted explaining today’s service issue.

There is currently an issue with one of the switches not responding in the cluster housed at bluesquare 2. An Engineer is being despatched to bluesquare to replace the faulty switch.

We apologise to any customer who may be affected during this time. We will update when more information becomes available.

Posted by Yolanda Zyczynska

Current Status

    GoodNo current issues

Status Archive